OpenStudy (anonymous):

What are solutions to the quadratic equation below? x2 - 4x + 4 = 3

OpenStudy (anonymous):

x= -b+-sqr b^2-4ac/2a

OpenStudy (anonymous):

start with \[x^2-4x+1=0\] and then use the quadratic formula, or else start with \[x^2-4x=-1\]and complete the square. in this case the second method is probably easiest

OpenStudy (anonymous):

apply this

OpenStudy (anonymous):

i will write out the second one, somone else will probably do the first \[x^2-4x=-1\] \[(x-2)^2=-1+2^2=3\] \[x-2=\pm\sqrt{3}\] \[x=2\pm\sqrt{3}\]

OpenStudy (anonymous):

whats the first one ?

OpenStudy (anonymous):

its a formula for finding the value for " x "

OpenStudy (anonymous):